Scores on the Scholastic Aptitude Test (SAT), a well known test taken by high school students are often converted to standardized scores or z-scores. The formula for determining a z-score on a mathematical portion of the SAT is given by z=(x-518)/115

where x is an individual score on the test. The value 518 is the national average score on the test, whereas 115 is a measure of how spread out the scores are.
1. Solve this formula for x in terms of z.
2. If your score is 2.2, what was your score on the test?

Answer by stanbon(48575)

The formula for determining a z-score on a mathematical portion of the SAT is given by z=(x-518)/115 where x is an individual score on the test. The value 518 is the national average score on the test, whereas 115 is a measure of how spread out the scores are.
1. Solve this formula for x in terms of z.
x-518 = 115z
x = 115z+518
------------------
2. If your z-score is 2.2, what was your x-score on the test?
x = 115*2.2 +518
x = 771
===============
Cheers,
Stan H.